OpenStudy (anonymous):

how to integrate (1/(x+1)^2)

OpenStudy (shubha):

(1/(x+1)^2) = (x+1)^(-2) now integral of y^n = (y^(n+1))/(n+1) so in this case y=x+1 and n=-2 when we integrate we get -1/(x+1)

OpenStudy (anonymous):

Have you covered integration by substitution yet?